# From Handcrafted Features to LLMs: A Brief Survey for Machine Translation Quality Estimation

### v1

@@ Abstract
As a critical subfield within NLP, MT has witnessed groundbreaking developments with the advent of deep learning technologies. Nonetheless, the quality of MT remains inherently uncertain. Traditional evaluation metrics, such as BLEU [papineni2002bleu], METEOR [banerjee2005meteor], TER [snover2006study], and CHRF rely on reference translations to assess translation quality. In contrast, QE techniques are capable of automatically evaluating the quality of translations without the need for reference, offering a valuable alternative for appraising the performance of MT systems.
@@
In the initial stages of QE for MT, there was no unified and clear definition of the field, and the research primarily concentrated on statistical machine translation systems. In 2009, researchers such as Specia et al. [specia2009estimating] introduced an innovative QE framework, which involved manual scoring annotations on translation, the implementation of feature engineering, and the use of ML algorithms to train models capable of predicting translation quality. Since the Workshop on Machine Translation (WMT) established QE as a separate task in 2012, the research has evolved into three main methods: the first is handcrafted feature-based QE; the second leverages deep learning for QE, which further includes subdivisions such as classic deep learning approaches and those incorporating pre-trained LMs; and the third, an emerging method, is based on LLMs. The development of these methods has significantly advanced the progress of QE and gradually improved the accuracy of QE models’ assessments.
Undoubtedly, approaches based on LLMs have become a focal point of research within the QE domain. Researchers are seeking to harness the extensive knowledge base and learning capabilities of LLMs to achieve new breakthroughs in QE studies. Currently, QE research based on LLMs mainly encompasses the following directions: first, direct prediction of translation quality scores [kocmi2023large], error levels [lu2023error], or fluency [yang2023knowledge] using LLMs; second, leveraging the generative probabilities from LLMs, which involves using various prompts and examples to obtain multiple generative probabilities for the translated sentences of source texts, thereby calculating the mean and variance to gain a more accurate measure of uncertainty for translation quality [huang2023towards]; third, generating pseudo data based on the knowledge within LLMs to then transfer to QE models [xu2023instructscore, HUANG2024102022]; fourth, employing LLMs as pre-trained foundation models to enhance QE systems [xu2023instructscore, gladkoff2023predicting]; and fifth, adopting retrieval-based approaches to infuse translation knowledge into LLMs [huang2023towards, HUANG2024102022]. Although the performance of LLM-based QE methods has not yet surpassed that of QE methods incorporating pre-trained LMs, it is anticipated that with ongoing research, LLM-based approaches have the potential to reach state-of-the-art (SOTA) performance levels.
@@
In this paper, our aim is to provide a clear and concise overview of MTQE for practitioners engaged in QE research and scholars interested in entering this field. In contrast to shared task overviews, our work not only synthesizes the WMT QE shared tasks from the past four years but also broadens the scope of content. Specifically, this paper reviews datasets, annotation methods, shared tasks, and all the seminal classic methods within the QE domain, with a particular emphasis on the currently highly-regarded QE approaches based on LLMs. Moreover, we explore the specific impact of LLMs on QE, a topic not yet covered in other survey reviews. Ultimately, we engage in an in-depth discussion of the current challenges faced by QE and the future research directions in this area.
@@ I-A Outline
In Section III, we review different methods in the QE field and classify them into feature engineering and ML-based methods, deep learning-based methods, and LLM-based methods. Among deep learning-based methods, we further categorize them into classic deep learning methods and those incorporating pre-trained LMs. We also list notable methods in each box of Fig. 1. In Section IV, we list five major challenges that currently exist in the QE field. Finally, in Section V, we provide our conclusions.
@@ II-A Datasets
The MLQE-PE dataset [fomicheva2020mlqepe] is a significant milestone in QE and Automatic Post-Editing (APE) research, providing annotations in a multilingual environment. The dataset is constructed using sentences from Wikipedia and Reddit articles. Parallel corpora are generated for 11 different language pairs (LPs), including 7 conventional resource LPs (English-German – En-De, English-Chinese – En-Zh, Russian-English – Ru-En, Romanian-English – Ro-En, Estonian-English – Et-En, Nepali-English – Ne-En, and Sinhala-English – Si-En) with 10K sentences each, divided into training, development, and two test sets (test20 and test21). Additionally, the dataset includes 4 zero-shot LPs (Pashto-English – Ps-En, Khmer-English – Km-En, English-Japanese – En-Ja, and English-Czech – En-Cs) with 2K sentences each, evenly split into two test sets too.
@@
These datasets provide an extremely important resource for research in the field of QE. They offer researchers corpora with rich texts and detailed annotations, which have propelled the progress of QE research. Specific related information can be found in Table I. It should be noted that, for the sake of brevity in presentation, the WMT2023 QE dataset omits all LPs from the MLQE-PE dataset.
| Dataset | LPs | Sentences | Tokens | Annotations | Data Source | Release Date | ||||||||||||||||
| Train | Dev | Test | Train | Dev | Test | DA | PE | MQM | ||||||||||||||
| MLQE-PE | En-De | 7,000 | 1,000 | 1,000/1,000 | 114,980 | 16,519 | 16,371/16,545 | ✓ | ✓ | Wikipedia | 2021/22 | |||||||||||
| En-Zh | 7,000 | 1,000 | 1,000/1,000 | 115,585 | 16,307 | 16,765/16,637 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Ru-En | 7,000 | 1,000 | 1,000/1,000 | 82,229 | 11,992 | 11,760/11,650 | ✓ | ✓ | 2021/22 | |||||||||||||
| Ro-En | 7,000 | 1,000 | 1,000/1,000 | 120,198 | 17,268 | 17,001/17,359 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Et-En | 7,000 | 1,000 | 1,000/1,000 | 98,080 | 14,423 | 14,358/14,044 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Ne-En | 7,000 | 1,000 | 1,000/1,000 | 104,934 | 15,144 | 14,770/15,017 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Si-En | 7,000 | 1,000 | 1,000/1,000 | 109,515 | 15,708 | 15,821/15,709 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Ps-En | - | 1,000 | 1,000 | - | 27,045 | 27,414 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| Km-En | - | 1,000 | 1,000 | - | 21,981 | 22,048 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| En-Ja | - | 1,000 | 1,000 | - | 20,626 | 20,646 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| En-Cs | - | 1,000 | 1,000 | - | 20,394 | 20,244 | ✓ | ✓ | Wikipedia | 2021/22 | ||||||||||||
| WMT2023 QE | En-Mr | 27,000 | 1,000 | 1,086 | 717,581 | 26,253 | 27,951 | ✓ | ✓ | multi-domain/multi-corpus | 2023 | En-Hi | 7,000 | 1,000 | 1,074 | 181,336 | 25,943 | 28,032 | ✓ | multi-domain/multi-corpus | 2023 | |
| En-Gu | 7,000 | 1,000 | 1,075 | 153,685 | 21,238 | 23,084 | ✓ | multi-domain/multi-corpus | 2023 | |||||||||||||
| En-Ta | 7,000 | 1,000 | 1,067 | 150,670 | 21,655 | 20,342 | ✓ | multi-domain/multi-corpus | 2023 | |||||||||||||
| En-Te | 7,000 | 1,028 | 1,000 | 147,492 | 20,686 | 22,640 | ✓ | multi-domain/multi-corpus | 2023 | |||||||||||||
| En-Fa | - | - | 1,000 | - | - | 26,807 | ✓ | news (multi-domain) | 2023 | |||||||||||||
| En-De | 30,425 | - | 1,897 | 877,066 | - | 37,996 | ✓ | multi-domain | 2021/23 | |||||||||||||
| En-Ru | 17,144 | - | - | 395,045 | - | - | ✓ | multi-domain | 2021/22 | |||||||||||||
| Zh-En | 36,851 | - | 1,675 | 1,654,454 | - | 39,770 | ✓ | multi-domain | 2021/23 | |||||||||||||
| He-En | - | - | 1,182 | - | - | 35,592 | ✓ | multi-domain | 2023 | |||||||||||||
@@ II-B1 Human Translation Error Rate (HTER)
HTER is a common method used for annotating translated sentences based on the volume of PE effort required. It builds upon word-level QE and is calculated from the results thereof. The goal of the reference translation is to make the least possible modifications while maintaining the original meaning and grammatical correctness. HTER scores a translated sentence by calculating the proportion of the number of edits (insertions, deletions, and replacements) made during the PE process to the number of words in the post-edition, with the formula provided in (1). In previous research, HTER has been analyzed as a substitute for human assessment, with some studies [Snover2005ASO] recommending its use as a gold standard for evaluation. However, opinions vary within the academic community regarding the adequacy of HTER as a substitute [Graham2016IsAT].
@@ II-B2 Direct Assessment (DA)
DA is a widely utilized manual evaluation method that provides a subjective quality assessment, taking into account the overall effect of translation outputs and serving as an alternative to HTER. During the DA evaluation process, annotators rate the quality of translations directly within a range of 0-100. When utilizing multiple DA scores as targets for QE tasks, these scores are typically normalized first, and the normalized averages are then used to represent the quality score of the MT output. DA scores are prone to inconsistency due to the influence of annotators’ individual preferences. However, solutions have been proposed to enhance the consistency among annotators [graham-etal-2013-continuous, guzman-etal-2019-flores]. As a result, DA has established itself as a reliable manual evaluation method [shapira2019crowdsourcing] and is extensively employed in QE tasks. Some advocate human assessments using DA, while others [fomicheva2020mlqepe] believe that DA and HTER offer distinct perspectives on MT quality. Both viewpoints are considered valid.
@@ II-B3 Multi-dimensional Quality Metrics (MQM)
MQM [lommel2014multidimensional] is an innovative and more objective annotation method that combines multiple evaluation indicators. It divides MT errors into 7 dimensions: Terminology, Accuracy, Linguistic conventions, Style, Locale conventions, Audience appropriateness, and Design and markup. Each dimension is subdivided into different error types, allowing for a more fine-grained QE of MT. Each dimension corresponds to four severity levels: no errors, minor errors, major errors, and critical errors, with different penalty scores for each level. Annotators can adjust parameters based on specific needs and integrate MQM into specific scenarios. Points are deducted based on the type of errors triggered by the MT, and the MT scores are calculated by subtracting the penalty scores from the full scores. MQM provides a more comprehensive and objective evaluation of MT quality compared to HTER and DA. It offers flexibility and personalized assessments, but requires annotators with domain knowledge and careful parameter settings. The formula is shown as in (2), where , , , and correspond to the counts of minor errors, major errors, critical errors, and the total number of words, respectively.
@@ II-C1 Word-level QE Shared Tasks
The primary evaluation metric for word-level QE is the Matthews correlation coefficient (MCC) [lin2004automatic], accompanied by the F1-score as a secondary metric. MCC is particularly suited for binary classification models and datasets with an uneven distribution. It is used to measure the correlation between incorrectly translated words and manual annotations.
@@ II-C2 Sentence-level QE Shared Tasks
Sentence-level QE aims to predict the quality score for each LP, indicative of the translation quality, akin to a regression task in ML. It employs annotation methods such as HTER, DA, and MQM [freitag2021experts] to assess the quality of translations. The sentence-level QE tasks at WMT2021222https://www.statmt.org/wmt21/quality-estimation-task.html, WMT2022, and WMT2023333https://wmt-qe-task.github.io/subtasks/task1/ have employed HTER, DA, and MQM annotations.
The primary evaluation metric for sentence-level QE is the Spearman’s [lin2004automatic, specia2009estimating] rank correlation coefficient (Spearman’s ), while the Pearson’s and Kendall’s correlation coefficients are used as auxiliary evaluation metrics. The Spearman’s does not rely on assumptions of normality and equal variance in translation quality scores and is less affected by outliers. Thus, it provides a better reflection of the correlation between the translation quality predicted by the MT model and manual annotation.
@@ III Methods of Quality Estimation
This section reviews relevant research work within the three main categories of methods that have emerged throughout the evolution of QE. It discusses the advantages and limitations of the respective methods within each category and provides a brief comparison between different approaches.
@@ III-A Quality Estimation Based on Handcrafted Features
Before 2009, QE research primarily focused on predicting quality labels for the output of Statistical Machine Translation (SMT) using handcrafted features [blatz2004confidence, ueffing2005word, ueffing2007word]. Subsequently, the focus of QE research shifted towards predicting human-annotated quality scores. For instance, the QuEst [specia-etal-2013-quest] framework utilized a feature extraction module to extract quality labels from the source and translated text. These features were then applied to ML algorithms to construct QE systems. de Souza et al. [de2014fbk] used a supervised tree-based ensemble learning method to predict PE effort and time under various features and BLSTM-RNNs to predict word-level labels.
QuEst++ [specia2015multi] is an improved and expanded version of QuEst, with added feature extraction modules designed for word-level and document-level QE. It integrates predictions at three different levels into a single workflow, facilitating interactions between word-level, sentence-level, and document-level QE. Additionally, QuEst++ incorporates sequence labeling learning algorithms for word-level QE. This tool can be conveniently extended with new features to meet the requirements of different text levels, offering high flexibility.

### v2

@@ Abstract
As a critical subfield within NLP, MT has witnessed groundbreaking developments with the advent of deep learning technologies. Nonetheless, the quality of MT remains inherently uncertain. Traditional evaluation metrics, such as BLEU [1], METEOR [2], TER [3], and CHRF rely on reference translations to assess translation quality. In contrast, QE techniques are capable of automatically evaluating the quality of translations without the need for reference, offering a valuable alternative for appraising the performance of MT systems.
@@
In the initial stages of QE for MT, there was no unified and clear definition of the field, and the research primarily concentrated on statistical machine translation systems. In 2009, researchers such as Specia et al. [4] introduced an innovative QE framework, which involved manual scoring annotations on translation, the implementation of feature engineering, and the use of ML algorithms to train models capable of predicting translation quality. Since the Workshop on Machine Translation (WMT) established QE as a separate task in 2012, the research has evolved into three main methods: the first is handcrafted feature-based QE; the second leverages deep learning for QE, which further includes subdivisions such as classic deep learning approaches and those incorporating pre-trained LMs; and the third, an emerging method, is based on LLMs. The development of these methods has significantly advanced the progress of QE and gradually improved the accuracy of QE models’ assessments. We have categorized these methods just to facilitate the description of the development trend of QE, and cannot guarantee there is no overlap between the methods in different categories.
Undoubtedly, approaches based on LLMs have become a focal point of research within the QE domain. Researchers are seeking to harness the extensive knowledge base and learning capabilities of LLMs to achieve new breakthroughs in QE studies. Currently, QE research based on LLMs mainly encompasses the following directions: first, direct prediction of translation quality scores [5], error levels [6], or fluency [7] using LLMs; second, leveraging the generative probabilities from LLMs, which involves using various prompts and examples to obtain multiple generative probabilities for the translated sentences of source texts, thereby calculating the mean and variance to gain a more accurate measure of uncertainty for translation quality [8]; third, generating pseudo data based on the knowledge within LLMs to then transfer to QE models [9, 10]; fourth, employing LLMs as pre-trained foundation models to enhance QE systems [9, 11]; and fifth, adopting retrieval-based approaches to infuse translation knowledge into LLMs [8, 10]. Although the performance of LLM-based QE methods has not yet surpassed that of QE methods incorporating pre-trained LMs, it is anticipated that with ongoing research, LLM-based approaches have the potential to reach state-of-the-art (SOTA) performance levels.
@@
In this paper, our aim is to provide a clear and concise overview of MTQE for practitioners engaged in QE research and scholars interested in entering this field. In contrast to shared task overviews, our work not only synthesizes the WMT QE shared tasks from the past four years but also broadens the scope of content. Specifically, this paper reviews datasets, annotation methods, shared tasks, and all the seminal classic methods within the QE domain, with a particular emphasis on the currently highly-regarded QE approaches based on LLMs. Moreover, we explore the specific impact of LLMs on QE, a topic not yet covered in other survey reviews. Ultimately, we engage in an in-depth discussion of the current challenges faced by QE and the future research directions.
@@ I-A Outline
In Section III, as shown in Fig. 1, we review different methods in the QE field and classify them into handcrafted features-based methods, deep learning-based methods, and LLM-based methods. Among deep learning-based methods, we further categorize them into classic deep learning methods and those incorporating pre-trained LMs. In Section IV, we list five major challenges that currently exist in the QE field. Finally, in Section V, we provide our conclusions.
@@ II-A Datasets
The MLQE-PE dataset [12] is a significant milestone in QE and Automatic Post-Editing (APE) research, providing annotations in a multilingual environment. The dataset is constructed using sentences from Wikipedia and Reddit articles. Parallel corpora are generated for 11 different language pairs (LPs), including 7 conventional resource LPs (English-German – En-De, English-Chinese – En-Zh, Russian-English – Ru-En, Romanian-English – Ro-En, Estonian-English – Et-En, Nepali-English – Ne-En, and Sinhala-English – Si-En) with 10K sentences each, divided into training, development, and two test sets (test20 and test21). Additionally, the dataset includes 4 zero-shot LPs (Pashto-English – Ps-En, Khmer-English – Km-En, English-Japanese – En-Ja, and English-Czech – En-Cs) with 2K sentences each, evenly split into two test sets too.
@@
These datasets provide an extremely important resource for research in the field of QE. They offer researchers corpora with rich texts and detailed annotations, which have propelled the progress of QE research. Specific related information can be found in Table LABEL:tab:dataset. It should be noted that, for the sake of brevity in presentation, the WMT2023 QE dataset omits all LPs from the MLQE-PE dataset.
Dataset | LPs | Sentences | Tokens | Annotations | Data Source | Release Date | ||||||
Train | Dev | Test | Train | Dev | Test | DA | PE | MQM | ||||
MLQE-PE | En-De | 7,000 | 1,000 | 1,000/1,000 | 114,980 | 16,519 | 16,371/16,545 | ✓ | ✓ | Wikipedia | 2021/22 | |
En-Zh | 7,000 | 1,000 | 1,000/1,000 | 115,585 | 16,307 | 16,765/16,637 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Ru-En | 7,000 | 1,000 | 1,000/1,000 | 82,229 | 11,992 | 11,760/11,650 | ✓ | ✓ | 2021/22 | |||
Ro-En | 7,000 | 1,000 | 1,000/1,000 | 120,198 | 17,268 | 17,001/17,359 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Et-En | 7,000 | 1,000 | 1,000/1,000 | 98,080 | 14,423 | 14,358/14,044 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Ne-En | 7,000 | 1,000 | 1,000/1,000 | 104,934 | 15,144 | 14,770/15,017 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Si-En | 7,000 | 1,000 | 1,000/1,000 | 109,515 | 15,708 | 15,821/15,709 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Ps-En | - | 1,000 | 1,000 | - | 27,045 | 27,414 | ✓ | ✓ | Wikipedia | 2021/22 | ||
Km-En | - | 1,000 | 1,000 | - | 21,981 | 22,048 | ✓ | ✓ | Wikipedia | 2021/22 | ||
En-Ja | - | 1,000 | 1,000 | - | 20,626 | 20,646 | ✓ | ✓ | Wikipedia | 2021/22 | ||
En-Cs | - | 1,000 | 1,000 | - | 20,394 | 20,244 | ✓ | ✓ | Wikipedia | 2021/22 | ||
WMT2023 QE | En-Mr | 27,000 | 1,000 | 1,086 | 717,581 | 26,253 | 27,951 | ✓ | ✓ | multi-domain/multi-corpus | 2023 | |
En-Hi | 7,000 | 1,000 | 1,074 | 181,336 | 25,943 | 28,032 | ✓ | multi-domain/multi-corpus | 2023 | |||
En-Gu | 7,000 | 1,000 | 1,075 | 153,685 | 21,238 | 23,084 | ✓ | multi-domain/multi-corpus | 2023 | |||
En-Ta | 7,000 | 1,000 | 1,067 | 150,670 | 21,655 | 20,342 | ✓ | multi-domain/multi-corpus | 2023 | |||
En-Te | 7,000 | 1,028 | 1,000 | 147,492 | 20,686 | 22,640 | ✓ | multi-domain/multi-corpus | 2023 | |||
En-Fa | - | - | 1,000 | - | - | 26,807 | ✓ | news (multi-domain) | 2023 | |||
En-De | 30,425 | - | 1,897 | 877,066 | - | 37,996 | ✓ | multi-domain | 2021/23 | |||
En-Ru | 17,144 | - | - | 395,045 | - | - | ✓ | multi-domain | 2021/22 | |||
Zh-En | 36,851 | - | 1,675 | 1,654,454 | - | 39,770 | ✓ | multi-domain | 2021/23 | |||
He-En | - | - | 1,182 | - | - | 35,592 | ✓ | multi-domain | 2023 |
@@ II-B1 Human Translation Error Rate (HTER)
HTER is a common method used for annotating translated sentences based on the volume of PE effort required. It builds upon word-level QE and is calculated from the results thereof. The goal of the reference translation is to make the least possible modifications while maintaining the original meaning and grammatical correctness. HTER scores a translated sentence by calculating the proportion of the number of edits (insertions, deletions, and replacements) made during the PE process to the number of words in the post-edition, with the formula provided in (1). In previous research, HTER has been analyzed as a substitute for human assessment, with some studies recommending its use as a gold standard for evaluation. However, opinions vary within the academic community regarding the adequacy of HTER as a substitute.
@@ II-B2 Direct Assessment (DA)
DA is a widely utilized manual evaluation method that provides a subjective quality assessment, taking into account the overall effect of translation outputs and serving as an alternative to HTER. During the DA evaluation process, annotators rate the quality of translations directly within a range of 0-100. When utilizing multiple DA scores as targets for QE tasks, these scores are typically normalized first, and the normalized averages are then used to represent the quality score of the MT output. DA scores are prone to inconsistency due to the influence of annotators’ individual preferences. However, solutions have been proposed to enhance the consistency among annotators [13]. As a result, DA has established itself as a reliable manual evaluation method and is extensively employed in QE tasks. Some advocate human assessments using DA, while others [12] believe that DA and HTER offer distinct perspectives on MT quality. Both viewpoints are considered valid.
@@ II-B3 Multi-dimensional Quality Metrics (MQM)
MQM [14] is an innovative and more objective annotation method that combines multiple evaluation indicators. It divides MT errors into 7 dimensions: Terminology, Accuracy, Linguistic conventions, Style, Locale conventions, Audience appropriateness, and Design and markup. Each dimension is subdivided into different error types, allowing for a more fine-grained QE of MT. Each dimension corresponds to four severity levels: no errors, minor errors, major errors, and critical errors, with different penalty scores for each level. Annotators can adjust parameters based on specific needs and integrate MQM into specific scenarios. Points are deducted based on the type of errors triggered by the MT, and the MT scores are calculated by subtracting the penalty scores from the full scores. MQM provides a more comprehensive and objective evaluation of MT quality compared to HTER and DA. It offers flexibility and personalized assessments, but requires annotators with domain knowledge and careful parameter settings. The formula is shown as in (2), where , , , and correspond to the counts of minor errors, major errors, critical errors, and the total number of words, respectively.
@@ II-C1 Word-level QE Shared Tasks
The primary evaluation metric for word-level QE is the Matthews correlation coefficient (MCC), accompanied by the F1-score as a secondary metric. MCC is particularly suited for binary classification models and datasets with an uneven distribution. It is used to measure the correlation between incorrectly translated words and manual annotations.
@@ II-C2 Sentence-level QE Shared Tasks
Sentence-level QE aims to predict the quality score for each LP, indicative of the translation quality, akin to a regression task in ML. It employs annotation methods such as HTER, DA, and MQM to assess the quality of translations. The sentence-level QE tasks at WMT2021222https://www.statmt.org/wmt21/quality-estimation-task.html, WMT2022, and WMT2023333https://wmt-qe-task.github.io/subtasks/task1/ have employed HTER, DA, and MQM annotations.
The primary evaluation metric for sentence-level QE is the Spearman’s rank correlation coefficient (Spearman’s ), while the Pearson’s and Kendall’s correlation coefficients are used as auxiliary evaluation metrics. The Spearman’s does not rely on assumptions of normality and equal variance in translation quality scores and is less affected by outliers. Thus, it provides a better reflection of the correlation between the translation quality predicted by the MT model and manual annotation.
@@ III Methods of Quality Estimation
This section reviews relevant research work within the three main categories of methods that have emerged throughout the evolution of QE. It discusses the advantages and limitations of the respective methods within each category and provides a brief comparison between different approaches. We have categorized these methods just to facilitate the description of the development trend of QE, and cannot guarantee there is no overlap between the methods in different categories.
@@ III-A Quality Estimation Based on Handcrafted Features
Before 2009, QE research primarily focused on predicting quality labels for the output of Statistical Machine Translation (SMT) using handcrafted features [15, 16]. Subsequently, the focus of QE research shifted towards predicting human-annotated quality scores. For instance, the QuEst [17] framework utilized a feature extraction module to extract quality labels from the source and translated text. These features were then applied to ML algorithms to construct QE systems. de Souza et al. [18] used a supervised tree-based ensemble learning method to predict PE effort and time under various features and BLSTM-RNNs to predict word-level labels.
QuEst++ [19] is an improved and expanded version of QuEst, with added feature extraction modules designed for word-level and document-level QE. It integrates predictions at three different levels into a single workflow, facilitating interactions between word-level, sentence-level, and document-level QE. Additionally, QuEst++ incorporates sequence labeling learning algorithms for word-level QE. This tool can be conveniently extended with new features to meet the requirements of different text levels, offering high flexibility.